# Training on the Benchmark Is Not All You Need

### v1

@@ Abstract
The success of Large Language Models (LLMs) relies heavily on the huge amount of pre-training data learned in the pre-training phase. The opacity of the pre-training process and the training data causes the results of many benchmark tests to become unreliable. If any model has been trained on a benchmark test set, it can seriously hinder the health of the field. In order to automate and efficiently test the capabilities of large language models, numerous mainstream benchmarks adopt a multiple-choice format. As the swapping of the contents of multiple-choice options does not affect the meaning of the question itself, we propose a simple and effective data leakage detection method based on this property. Specifically, we shuffle the contents of the options in the data to generate the corresponding derived data sets, and then detect data leakage based on the model’s log probability distribution over the derived data sets. If there is a maximum and outlier in the set of log probabilities, it indicates that the data is leaked. Our method is able to work under black-box conditions without access to model training data or weights, effectively identifying data leakage from benchmark test sets in model pre-training data, including both normal scenarios and complex scenarios where options may have been shuffled intentionally or unintentionally. Through experiments based on two LLMs and benchmark designs, we demonstrate the effectiveness of our method. In addition, we evaluate the degree of data leakage of 31 mainstream open-source LLMs on four benchmark datasets and give a ranking of the leaked LLMs for each benchmark, and we find that the Qwen family of LLMs has the highest degree of data leakage [^1] Code open source at: https://github.com/nishiwen1214/Benchmark-leakage-detection.
@@ Introduction
In this work, we show how to provide reliable evidence for test set contamination in black-box language models. More specifically, we provide a simple and efficient new method for benchmark leakage detection based on multiple-choice questions. The method identifies the presence of a benchmark test set in a language model’s pre-training data and the extent of data leakage without accessing the model’s training data or weights. The contributions of this paper are summarized below:
@@
  • The algorithms are able to work in black-box conditions without access to model training data or weights, effectively identifying data leakage from the benchmark test set in the model pre-training data, including normal scenarios and complex scenarios in which options may have been intentionally or unintentionally disrupted.
  • We validate the effectiveness of the approach based on two LLMs design experiments and evaluate the data leakage risk of 31 open-source LLMs on four mainstream benchmark sets, present a benchmark leakage leaderboard among LLMs, and in particular find that the Qwen family of LLMs shows a high risk in several benchmarks.
@@
Our work focuses on the problem of data leakage of LLMs on benchmark test sets. We therefore discuss the work most relevant to us from the perspectives of both mainstream large language model benchmarks and data leakage detection.
@@
The current pre-training model size and its pre-training corpus are getting larger and larger, which inevitably leads to data leakage between the pre-training corpus and various benchmark test sets. Several previous studies (Brown et al. 2020; Wei et al. 2021) have utilized post-hoc n-gram overlap analysis between the benchmark and pre-training corpus to measure data leakage. Deng et al. (2023) utilized benchmark perturbations and synthetic data to detect benchmark leakage. (Wei et al. 2023) compare the model’s loss on the training, validation, and test sets; if the model’s loss on the training set is significantly lower than on the validation or test sets, this may indicate that the model is overfitting the training data. If the loss on the test set is significantly lower than an independent reference set (consisting of data that the model has never seen), this may indicate that the test data was compromised during training. Mattern et al. (2023) tested the difference in perplexity between the target sequence and the randomized sequence. Oren et al. (2023) exchanged the order of problems in some benchmarks and tested the model with generating new data as a way to detect data leakage. Xu et al. (2024) introduced a measure of the predictive accuracy of the benchmark model using two simple and scalable metrics, complexity and n-gram accuracy, to identify potential data leakage. Our work leverages the interchangeability of options in benchmark test sets to achieve instance and fine-grained data leakage detection. In addition to common scenarios, we even consider data leakage identification in scenarios where options are intentionally or unintentionally shuffled.
@@
Our goal is to identify whether the pre-training process of a language model $\theta$ includes a particular piece of data $x$ from a benchmark test set, or the extent to which that benchmark test set $D$ leaks to the model $\theta$. Detection in our setup is under black-box conditions, i.e., the pre-training corpus and parameters of the model are unknown. We consider two scenarios: (a) where the order in which the pre-trained data options are presented is not shuffled, and (b) where the sequence of pre-trained data options may be shuffled.
@@
Table 1: Experiment results on MMLU and CMMLU datasets under scenario (a).
@@
In this section we experimentally demonstrate the effectiveness of our proposed method for detecting data leakage.
@@
The previous experiments demonstrate the effectiveness of our Algorithm 1 and Algorithm 2, and next we will construct leaderboards for various benchmark leaks of LLMs. We conduct comprehensive data leakage detection experiments on four mainstream benchmarks: MMLU (Hendrycks et al. 2021a), CMMLU (Li et al. 2023), C-Eval (Huang et al. 2024), CMB (Wang et al. 2023). As shown in Figuu’rere 3, we tested almost all of the currently popular 31 LLMs (Team 2023; Touvron et al. 2023a, b; Young et al. 2024; BAAI 2023; Bai et al. 2023; Yang et al. 2023; Bi et al. 2024; Abdin et al. 2024), and we give the percentage predicted to be data leakage for both scenarios a and b. The outlier threshold $\delta$ for our scenario b is set to 0.2 on the three benchmark test sets, MMLU, CMMLU, and C-Eval; since there are five options for the data in the CMB benchmark, its outlier threshold $\delta$ is set to 0.25. And Ordered by the degree of leakage under scenario b. The Benchmark leakage leaderboard in Figure 3 is sorted by the degree of leakage under the scenario b. First of all, we find that there is not much gap between models on the MMLU benchmark, and the top five models in terms of data leakage risk are Qwen2-72B, Qwen1.5-110B, Yi-34B, Yi1.5-9B and Yi1.5-6B. Overall, the leakage of LLMs on the MMLU benchmark is a serious concern, and as MMLU is one of the most used and widely used benchmarks in the English language domain, the issue deserves our attention.
@@
Overall, GLM4-9B has the lowest risk of data leakage on all three benchmarks, MMLU, CMMLU, and C-Eval, and a low risk of data leakage on CMB. Qwen family LLMs have very high leakage risk on all 4 benchmarks, and we find that the larger the model the higher the leakage index, which might be due to the fact that larger models have more pre-training data and are more capable of learning and remembering the data more firmly. In addition to the Qwen family LLMs, the Yi family LLMs, DeepSeek family LLMs, and Baichuan family LLMs are also at slight risk of benchmark compromise. Mild benchmark leaks are hard to avoid, but we hope that researchers should avoid serious benchmark leaks when developing LLMs.
@@ Conclusion
This work has highlighted the severity of benchmark data leakage in Large Language Models (LLMs) and introduced an innovative detection method capable of identifying leakages under various scenarios, including when the order of multiple-choice options may have been shuffled. We validate the effectiveness of the approach based on two LLMs design experiments and evaluate the data leakage risk of 31 open-source LLMs on four mainstream benchmark sets, present a benchmark leakage leaderboard among LLMs, and in particular find that the Qwen family of LLMs shows a high risk in several benchmarks. This work emphasizes the need for developers and researchers to be vigilant in ensuring the integrity and fairness of LLM assessments. We call for continued community effort to address this issue, improve our detection techniques, and uphold the robustness of benchmark assessments in the field of artificial intelligence. This paper serves as a stepping stone towards establishing more reliable and trustworthy standards in the evaluation of LLMs and advancing the field of artificial intelligence with confidence and integrity. Currently our method is limited to detecting data in multiple choice format, in the future we will try to extend our method to other formats. In addition some multimodal benchmarks are also in the format of multiple choice questions, and in the future we will also attempt to detect benchmark leakage on large multimodal models.

### v2

@@ Abstract
The success of Large Language Models (LLMs) relies heavily on the huge amount of pre-training data learned in the pre-training phase. The opacity of the pre-training process and the training data causes the results of many benchmark tests to become unreliable. If any model has been trained on a benchmark test set, it can seriously hinder the health of the field. In order to automate and efficiently test the capabilities of large language models, numerous mainstream benchmarks adopt a multiple-choice format. As the swapping of the contents of multiple-choice options does not affect the meaning of the question itself, we propose a simple and effective data leakage detection method based on this property. Specifically, we shuffle the contents of the options in the data to generate the corresponding derived data sets, and then detect data leakage based on the model’s log probability distribution over the derived data sets. If there is a maximum and outlier in the set of log probabilities, it indicates that the data is leaked. Our method is able to work under gray-box conditions without access to model training data or weights, effectively identifying data leakage from benchmark test sets in model pre-training data, including both normal scenarios and complex scenarios where options may have been shuffled intentionally or unintentionally. Through experiments based on two LLMs and benchmark designs, we demonstrate the effectiveness of our method. In addition, we evaluate the degree of data leakage of 35 mainstream open-source LLMs on four benchmark datasets and give a ranking of the leaked LLMs for each benchmark, and we find that the Qwen family of LLMs has the highest degree of data leakage.
Codehttps://github.com/nishiwen1214/Benchmark-leakage-detection
@@ Introduction
In this work, we show how to provide reliable evidence for test set contamination in gray-box language models (Only need to access the Log-Probability of the output). More specifically, we provide a simple and efficient new method for benchmark leakage detection based on multiple-choice questions. The method identifies the presence of a benchmark test set in a language model’s pre-training data and the extent of data leakage without accessing the model’s training data or weights. The contributions of this paper are summarized below:
@@
  • The algorithms are able to work in gray-box conditions without access to model training data or weights, effectively identifying data leakage from the benchmark test set in the model pre-training data, including normal scenarios and complex scenarios in which options may have been intentionally or unintentionally disrupted.
  • We validate the effectiveness of the approach based on two LLMs design experiments and evaluate the data leakage risk of 35 open-source LLMs on four mainstream benchmark sets, present a benchmark leakage leaderboard among LLMs, and in particular find that the Qwen family of LLMs shows a high risk in several benchmarks.
@@
The current pre-training model size and its pre-training corpus are getting larger and larger, which inevitably leads to data leakage between the pre-training corpus and various benchmark test sets. Several previous studies (Brown et al. 2020; Wei et al. 2021) have utilized post-hoc n-gram overlap analysis between the benchmark and pre-training corpus to measure data leakage. Deng et al. (2023) utilized benchmark perturbations and synthetic data to detect benchmark leakage. (Wei et al. 2023) compare the model’s loss on the training, validation, and test sets; if the model’s loss on the training set is significantly lower than on the validation or test sets, this may indicate that the model is overfitting the training data. If the loss on the test set is significantly lower than an independent reference set (consisting of data that the model has never seen), this may indicate that the test data was compromised during training. Mattern et al. (2023) tested the difference in perplexity between the target sequence and the randomized sequence. Oren et al. (2023) exchanged the order of problems in some benchmarks and tested the model with generating new data as a way to detect data leakage. Xu et al. (2024) introduced a measure of the predictive accuracy of the benchmark model using two simple and scalable metrics, complexity and n-gram accuracy, to identify potential data leakage. Dong et al. (2024) identified data contamination by analysing the peakedness of the model output distribution. Our work leverages the interchangeability of options in benchmark test sets to achieve instance and fine-grained data leakage detection. In addition to common scenarios, we even consider data leakage identification in scenarios where options are intentionally or unintentionally shuffled.
@@
Our goal is to identify whether the pre-training process of a language model $\theta$ includes a particular piece of data $x$ from a benchmark test set, or the extent to which that benchmark test set $D$ leaks to the model $\theta$. Detection in our setup is under gray-box conditions, i.e., the pre-training corpus and parameters of the model are unknown. We consider two scenarios: (a) where the order in which the pre-trained data options are presented is not shuffled, and (b) where the sequence of pre-trained data options may be shuffled.
@@
Table 1: Experiment results under scenario (a).
@@
The previous experiments demonstrate the effectiveness of our Algorithm 1 and Algorithm 2, and next we will construct leaderboards for various benchmark leaks of LLMs. We conduct comprehensive data leakage detection experiments on four mainstream benchmarks: MMLU (Hendrycks et al. 2021a), CMMLU (Li et al. 2023), C-Eval (Huang et al. 2024), CMB (Wang et al. 2023). As shown in Figure 3, we tested almost all of the currently popular 35 LLMs (Team 2023; Touvron et al. 2023a, b; Young et al. 2024; BAAI 2023; Bai et al. 2023; Yang et al. 2023; Bi et al. 2024; Abdin et al. 2024), and we give the percentage predicted to be data leakage for both scenarios a and b. The outlier threshold $\delta$ for our scenario b is set to 0.2 on the three benchmark test sets, MMLU, CMMLU, and C-Eval; since there are five options for the data in the CMB benchmark, its outlier threshold $\delta$ is set to 0.25. And Ordered by the degree of leakage under scenario b. The Benchmark leakage leaderboard in Figure 3 is sorted by the degree of leakage under the scenario b. First of all, we find that there is not much gap between models on the MMLU benchmark, and the top five models in terms of data leakage risk are Qwen2-72B, Qwen1.5-110B, Yi-34B, Yi1.5-9B and Yi1.5-6B. Overall, the leakage of LLMs on the MMLU benchmark is a serious concern, and as MMLU is one of the most used and widely used benchmarks in the English language domain, the issue deserves our attention.
@@
Overall, GLM4-9B and MiniCPM3-4B have the lowest risk of data leakage on all three benchmarks, MMLU, CMMLU, and C-Eval, and a low risk of data leakage on CMB. Qwen family LLMs have very high leakage risk on all 4 benchmarks, and we find that the larger the model the higher the leakage index, which might be due to the fact that larger models have more pre-training data and are more capable of learning and remembering the data more firmly. In addition to the Qwen family LLMs, the Yi family LLMs, DeepSeek family LLMs, and Baichuan family LLMs are also at slight risk of benchmark compromise. Mild benchmark leaks are hard to avoid, but we hope that researchers should avoid serious benchmark leaks when developing LLMs.
@@ Conclusion
This work has highlighted the severity of benchmark data leakage in Large Language Models (LLMs) and introduced an innovative detection method capable of identifying leakages under various scenarios, including when the order of multiple-choice options may have been shuffled. We validate the effectiveness of the approach based on two LLMs design experiments and evaluate the data leakage risk of 35 open-source LLMs on four mainstream benchmark sets, present a benchmark leakage leaderboard among LLMs, and in particular find that the Qwen family of LLMs shows a high risk in several benchmarks. This work emphasizes the need for developers and researchers to be vigilant in ensuring the integrity and fairness of LLM assessments. We call for continued community effort to address this issue, improve our detection techniques, and uphold the robustness of benchmark assessments in the field of artificial intelligence. This paper serves as a stepping stone towards establishing more reliable and trustworthy standards in the evaluation of LLMs and advancing the field of artificial intelligence with confidence and integrity. Currently our method is limited to detecting data in multiple choice format, in the future we will try to extend our method to other formats.
Acknowledgements
This work was supported by Open Research Fund of Key Laboratory of Intelligent Education Technology and Application of Zhejiang Province, Zhejiang Normal University (Grant No. jykf21002w), National Natural Science Foundation of China (62376262), GuangDong Basic and Applied Basic Research Foundation (2023A1515110718 and 2024A1515012003), China Postdoctoral Science Foundation (2024M753398), Postdoctoral Fellowship Program of CPSF (GZC20232873).